# Overcoming the Zero-Rate Hashing Bound with Holographic Quantum Error Correction

### v2

@@ IV Discussion
In this work we have shown strong evidence of extremely high threshold behavior for zero-rate holographic codes in the presence of 1-Pauli and 2-Pauli noise channels, for pure and finitely-biased noise regimes. In doing so, we have tested the: HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes; Steane code; $\llbracket 6,1,3\rrbracket$ code; SCF code; and a Clifford-deformed Steane code. All of these codes admit thresholds for some Pauli noise bias which either attain or surpass the hashing bound, and one of these codes, the SCF, surpasses known records for 2-Pauli noise. Our work thus demonstrates that holographic codes exhibit remarkable code-capacity properties under biased noise channels and can even overtake current state-of-the-art results [27]. It moreover indicates that they constitute a novel and competitive class of stabilizer codes that is robust against biased noise, beyond the conventional options of topological codes [25, 22, 26, 23, 21, 20], naïve constructions of concatenated codes [80], or advanced code-concatenation techniques [27]. Our results heavily imply that further modification via Clifford deformations is possible, and additionally that holographic codes can in fact be tailored to Pauli noise channels of interest.
Our work leaves many open questions for the future. Firstly, we mentioned previously that the limit of 1-Pauli noise (that is, $\eta\mapsto\infty$) for both the HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes achieve the zero-rate hashing bound. The results that we have shown are in line with those of [24, 38], wherein such basic seed tensors were built up as in the structure of cyclic XZZX codes, and then mapped to generalized toric codes. Here, although the threshold values that we find are perfectly in line with the results of [24], our approach differs substantially in that we have uncovered a new class of subsystem stabilizer code for which several variants can achieve the zero-rate hashing bound for 1-Pauli noise, in line with several other known code constructions [27, 20, 25, 22, 23, 21].
@@
In related work [56], an integer-optimization decoder was utilized for the zero-rate Evenbly code in order to pinpoint depolarizing and pure 1-Pauli noise thresholds of $19.1\%$ and $50\%$, respectively overcoming and achieving the hashing bound. It would be interesting to see if a tensor-network decoder would allow for a more complete biased-noise threshold exploration of Evenbly codes and their many derivatives [79, 81, 82]. However, in order to realize such an aim, the current tensor-network decoder utilized in our study would need to incorporate the extra Hadamard gates on edges, in order to extract the correct destabilizers from a syndrome. We leave such pursuits for subsequent work.
@@
Additionally, it is known from [56, 79] that Evenbly codes exhibit gauge-dependent threshold behavior, owing to the specific isometric constraints used in the structure of the code. In principle, nothing restricts us from utilizing such gauge-fixing techniques in conjunction with maximum-rate constructions of the holographic codes studied here. However, the unique seed-tensor isometry properties of each holographic code may be utilized in order to give rise to similar gauge-dependent threshold “phases”, as the maximum-rate HaPPY code under gauge-fixing can exhibit very high threshold behavior [83] on a $\{5,5\}$ hyperbolic tiling.
@@
Concerning the topic of finite-rate holographic code constructions, it is currently unknown whether there exist logical-qubit implantation schemes which maximize a code-capacity threshold for a given error model. It was shown in [37] that several naive logical-qubit implantation schemes can approach the hashing bound; furthermore, it has been conjectured in [56] that certain finite-rate conceptions of Evenbly codes may in fact achieve the code-capacity threshold for quantum erasure noise, as well as the finite-rate hashing bound for Pauli noise. Investigating these possibilities will be at the center of future research efforts.
@@
Recently, generalizations of Abelian stabilizer codes have been investigated as alternatives for low-overhead quantum error correction, as such codes naturally allow for non-Clifford logical gates [50, 84, 85]. The quantum LEGO formalism itself is not restricted to only stabilizer circuits, and may be readily utilized for simulating aspects of either non-Abelian stabilizer or non-stabilizer codes. In this case, an enumerator-based method that combines tensor network and Monte Carlo sampling can be used to estimate logical error probabilities and thresholds of the code [41]. Alternatively, it is possible to calculate the coherent information instead [86, 87]. However, new techniques are likely needed to calculate the coherent information for such generalized holographic codes, since the number of physical qubits at the boundary grows exponentially with a commensurate number of layers [35, 54].
@@
Comparing all of our results with those of concatenated codes, we note that the codes tested here fare better than typical tree-style concatenation [80, 88], particularly for the depolarizing noise channel. Nevertheless, it was shown in [27] that different types of repetition code concatenations can achieve threshold values comparable to ours for the depolarizing noise limit. Additionally, it was shown that various different concatenation schemes achieve or slightly surpass the hashing bound, depending on the details of the two repetition codes involved in the concatenation. Given this, one may then ask as to why our zero-rate holographic code constructions behave as they do. Considering the structural differences between holographic codes and more straightforward code-concatenation schemes, we would expect that in the 2-Pauli limit a conferrable advantage must be present by concatenation along the edges of a hyperbolic tessellation. Our forthcoming work will investigate this advantage more thoroughly.
@@
As a final observation from the more theoretical perspective, recent work [65] has shown that the bulk geometric transition in AdS/CFT implies a threshold for holographic codes in the continuum limit. Taking this finding in tandem with our results, it is natural to ask whether the discretized isometric map typical of holographic codes provides a general method for achieving above-hashing-bound behavior, and additionally, whether or not AdS/CFT in the continuum limit implies the existence of a fault-tolerance threshold for holographic codes.
IV.1 Holographic Codes in Practice
@@
In this work, we have investigated the threshold of holographic codes with respect to code-capacity noise channels. Although it has not been confirmed whether or not holographic codes admit a fault-tolerance threshold [65], it seems unreasonable to suspect otherwise, given the fact that the threshold data we have presented here is comparable to many state-of-the-art code families. A first step could be to test phenomenological noise models, i.e., those that incorporate measurement noise into the syndrome-extraction process. This goal could in principle be achieved by mapping the tensor network to a detector picture, as was recently performed in [89], thereby facilitating large-scale circuit-level noise simulations.
@@
In the interim, one can consider techniques by which full fault-tolerant syndrome extraction and universal logic can be performed. [90, 91, 92] proved that generalized concatenated code schemes based on the quantum Hamming code can exhibit constant space overhead and quasi-polylogarithmic time overhead. These schemes, all based on code concatenation, involve the use of hard-decision layerwise decoding protocols [93]. As it has been proven that holographic quantum codes are specific instances of generalized concatenated codes for the hyperbolic plane [67], we would surmise that the aforementioned techniques can also be adapted to our setting. This idea is the subject of active research.
@@
Many alternatives have been suggested as well to the layerwise decoding scheme. One method may involve floquetifying holographic codes [94]; this procedure would solve two problems at once, in that high-weight stabilizers and logical operators could be controlled to a reasonable weight, and fault-tolerance can be guaranteed with an appropriate measurement schedule. In addition to these ideas, fault-tolerant logic has also been shown to be gauge-dependent [56] and universal in some cases for holographic codes [67]. Recent work has also made progress on lowering the computational complexity of decoding itself using tensor-network schemes [95, 96]. These techniques, often based on the analysis of critical or approximate contraction paths, could allow for even faster decoding than is currently employed by polynomial-complexity tensor-network decoding [37, 36], and may become a prime factor in the construction practical quantum codes which supersede the hashing bound.
@@
One may also ask what specific applications may be well-suited to the usage of a holographic code. An application of holographic codes may be found in magic-state distillation [97, 98]. Indeed, for the case of the HaPPY and holographic Steane codes, the transversal logical $\overline{SH}$ and $\bar{H}$ gates allow for current state-of-the-art methods to be leveraged, as was pointed out in [99]. For practical quantum error correction, it is known that the square-lattice GKP code admits an induced $XZ$ 2-Pauli noise channel under a symmetric Gaussian random displacement noise model [74, 75, 76, 77], and also that Pauli-twirling the amplitude damping channel results in biased 2-Pauli noise [78]. Such applications could be combined with our work in the future.
IV.2 Acknowledgements
@@
We thank Aritra Sarkar, David Elkouss, and Jens Eisert for discussions related to concatenated codes and quantum channel capacity. We also thank Mackenzie Shaw and Kathleen Chang for mentioning at the conference FTQC ’24 in Benasque several realistic examples of 2-Pauli noise arising in practical quantum computing. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation”.
IV.3 Author Contributions

### v3

@@ IV Discussion
In this work we have shown strong evidence of extremely high threshold behavior for zero-rate holographic codes in the presence of 1-Pauli and 2-Pauli noise channels, for pure and finitely-biased noise regimes. In doing so, we have tested the: HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes; Steane code; $\llbracket 6,1,3\rrbracket$ code; SCF code; and a Clifford-deformed Steane code. All codes have been shown to admit thresholds for some Pauli noise bias which either attain or surpass the hashing bound, and two of these codes, the SCF and the $\llbracket 6,1,3\rrbracket$ code, surpass known the current state of the art for 2-Pauli noise [leditzky_pauli_hashing, bhalerao2025improving]. Our work demonstrates that holographic codes exhibit remarkable code-capacity properties under biased noise channels and can even overtake current state-of-the-art results [leditzky_pauli_hashing]. It moreover indicates that they constitute a novel and competitive class of stabilizer codes that is robust against biased noise, beyond the conventional options of topological codes [xzzx_surface_code, basudha_xyz2, domain_wall_color_jens, tuckett_biased_surface, 3d_surface_hashing, clifford_deformed_surface], naïve constructions of concatenated codes [exact_performance], or other advanced techniques [leditzky_pauli_hashing]. Our results heavily imply that further modification via Clifford deformations is possible, and additionally that holographic codes can in fact be tailored to Pauli noise channels of interest.
@@
Although our work constitutes the first such study of biased noise and holographic quantum codes, there are still many open directions for future research. Firstly, In related work [evenbly_codes2], an integer-optimization decoder was utilized for the zero-rate Evenbly code in order to pinpoint depolarizing and pure 1-Pauli noise thresholds of $19.1\%$ and $50\%$, respectively overcoming and achieving the hashing bound. It would be interesting to see if a tensor-network decoder would allow for a more complete biased-noise threshold exploration of Evenbly codes and their many derivatives [steinberg_evenbly_codes1, evenblyHTN, steinberg_conformal_HTN]. However, in order to realize such an aim, the current tensor-network decoder utilized in our study would need to incorporate the extra Hadamard gates on edges, in order to extract the correct destabilizers from a syndrome. We leave such pursuits for subsequent work.
@@
Additionally, it is known from [evenbly_codes2, steinberg_evenbly_codes1] that Evenbly codes exhibit gauge-dependent threshold behavior, owing to the specific isometric constraints used in the structure of the code. In principle, nothing restricts us from utilizing such gauge-fixing techniques in conjunction with maximum-rate constructions of the holographic codes studied here. However, the unique seed-tensor isometry properties of each holographic code may be utilized in order to give rise to similar gauge-dependent threshold “phases”; one example of this can be seen in the maximum-rate HaPPY code under gauge-fixing, which can exhibit very high threshold behavior [alex_logical_recovery] on a $\{5,5\}$ hyperbolic tiling. A systematic exploration of this subtopic would help to clarify the behavior of holographic codes under gauge-fixing constraints.
@@
Thirdly, regarding the topic of finite-rate holographic code constructions, it is currently unknown whether there exist logical-qubit implantation schemes which maximize a code-capacity threshold for a given error model in a holographic code. It was shown in [parallel_decoding_tn_codes] that several naive logical-qubit implantation schemes can approach the hashing bound; furthermore, it has been conjectured in [evenbly_codes2] that certain finite-rate conceptions of Evenbly codes may in fact achieve the code-capacity threshold for quantum erasure noise, as well as the finite-rate hashing bound for Pauli noise. Investigating these possibilities will be at the center of future research efforts.
@@
Recently, generalizations of Abelian stabilizer codes have been investigated as alternatives for low-overhead quantum error correction, as such codes naturally allow for non-Clifford logical gates [xp_stabilizer, webster2022xp, turaev_viro]. The quantum LEGO formalism itself is not restricted to only stabilizer circuits, and may be readily utilized for simulating aspects of either non-Abelian stabilizer or non-stabilizer codes. In this case, an enumerator-based method that combines tensor network and Monte Carlo sampling can be used to estimate logical error probabilities and thresholds of the code [cao2024quantum]. Alternatively, it is possible to calculate the coherent information instead [approx_CFT_codes, coherent_information]. However, new techniques are likely needed to calculate the coherent information for such generalized holographic codes, since the number of physical qubits at the boundary grows exponentially with a commensurate number of layers [holographic_codes_topical_review, conformal_quasi_holo].
@@
Comparing all of our results with those of concatenated codes, we note that the codes tested here fare better than typical tree-style concatenation [exact_performance, chamberland2017complementarygauge], particularly for the depolarizing noise channel. Nevertheless, it was shown in [leditzky_pauli_hashing] that different types of repetition code concatenations can achieve threshold values comparable to ours for the depolarizing noise limit. Additionally, it was shown that various different concatenation schemes achieve or slightly surpass the hashing bound, depending on the details of the two repetition codes involved in the concatenation. Given this, one may then ask as to why our zero-rate holographic code constructions behave as they do. Considering the structural differences between holographic codes and more straightforward code-concatenation schemes, we would expect that in the 2-Pauli limit a conferrable advantage must be present by concatenation along the edges of a hyperbolic tessellation. Forthcoming work will investigate this advantage on more general footing.
@@
As a final observation from high-energy physics, recent work [deconfinement] has shown that the bulk geometric transition in AdS/CFT implies a threshold for holographic codes in the continuum limit. Taking this finding in tandem with our results, it is natural to ask whether the discretized isometric map typical of holographic codes provides a general method for achieving above-hashing-bound behavior, and additionally, whether or not AdS/CFT in the continuum limit implies the existence of a fault-tolerance threshold for holographic codes.
IV.1 Quantum Coding Theory
Here we put our results into perspective for quantum channel capacity and coding theory. Currently, understanding and mitigating the impact of quantum noise is crucial for quantum computation and communication. The most common way to characterize quantum noise is through quantum channels, which provide rigorous descriptions for how a quantum state is transformed or transmitted in the presence of errors [wilde_qit]; as such, the maximum number of qubits that can be theoretically transmitted reliably in a noisy environment is the quantum channel capacity (QCC) [bennett1998quantum, qccapacity_very_noisy, schumacher1996quantum, schumacher1996sending, privateclassical_qcapacity, lloyd_capacity_noisy]. Determining the precise value of the QCC is therefore known to be incredibly challenging [Bennett_1997], and in particular, the QCC is not accurately known even for a channel as ubiquitous as the depolarizing channel. A large part of the difficulty lies in the superadditivity of coherent information, which hinders a direct efficient evaluation of the quantum capacity when the channel acts independently and identically over many qubits. As a result, finding lower and upper bounds of the quantum capacity constitute important progress in quantum Shannon theory [wilde_qit, bennett1998quantum].
In this vein, our results stand alongside recent efforts to characterize new classes of quantum codes that come arbitrarily close to the QCC [leditzky_pauli_hashing, xzzx_surface_code]. As the leading method for generating practical QEC codes achieving the QCC relies on random stabilizer codes, our work demonstrates via the holographic code construction that generalizations of code concatenation can achieve the theoretical lower bound for the QCC (i.e., the hashing limit). Such results are of paramount interest, since, as stated above, random quantum stabilizer codes do not generally achieve the optimal transmission rate for their analogous quantum counterparts such as the depolarizing channel or other Pauli channels, unlike their classical counterparts (which are typically capacity-achieving for symmetric channels).
IV.2 Holographic Codes in Practice
@@
We close with a short discussion on the practical implications of our findings. In this work, we have investigated the threshold of holographic codes with respect to code-capacity noise channels. Although it has not been confirmed whether or not holographic codes admit a fault-tolerance threshold [deconfinement], it seems unreasonable to suspect otherwise, given the fact that the threshold data we have presented here is comparable to many state-of-the-art code families. A first step could be to test phenomenological noise models, i.e., those that incorporate measurement noise into the syndrome-extraction process. This goal could in principle be achieved by mapping the tensor network to a detector picture, as was recently performed in [renes_2d], thereby facilitating large-scale circuit-level noise simulations.
@@
In the interim, one can consider techniques by which full fault-tolerant syndrome extraction and universal logic can be performed. [yamasaki1, yamasaki2, goto_many_hypercube] proved that generalized concatenated code schemes based on the quantum Hamming code can exhibit constant space overhead and quasi-polylogarithmic time overhead. These schemes, all based on code concatenation, involve the use of hard-decision layerwise decoding protocols [knill2005quantum]. As it has been proven that holographic quantum codes are specific instances of generalized concatenated codes for the hyperbolic plane [heterogeneous_holo_qrm], we would surmise that the aforementioned techniques can also be adapted to our setting. This idea is the subject of ongoing research.
@@
Moreover, many alternatives have been suggested as well to the layerwise decoding scheme mentioned above. One method may involve floquetifying holographic codes [hastings2021dynamically]; this procedure would solve two problems at once, in that high-weight stabilizers and logical operators could be controlled to a reasonable weight, and fault-tolerance could be guaranteed with an appropriate measurement schedule. In addition to these ideas, fault-tolerant logic has also been shown to be gauge-dependent [evenbly_codes2] and universal in some cases for holographic codes [heterogeneous_holo_qrm]. Recent work has also made progress on lowering the computational complexity of decoding itself using tensor-network schemes [gray2021hyperkourtis, hyperoptimized_approximate]. These techniques, often based on the analysis of critical or approximate contraction paths, could allow for even faster decoding for error correction and logical operations than is currently employed by tensor-network decoding schemes [parallel_decoding_tn_codes, farrelly_tn_codes].
@@
One may also ask what specific applications may be well-suited to the usage of a holographic code. An application of holographic codes may be found in magic-state distillation [magic1, magic2]. Indeed, for the case of the HaPPY and holographic Steane codes, the transversal logical $\overline{SH}$ and $\bar{H}$ gates allow for current state-of-the-art methods to be leveraged, as was pointed out in [magic3]. For practical quantum error correction, it is known that the square-lattice GKP code admits an induced $XZ$ 2-Pauli noise channel under a symmetric Gaussian random displacement noise model [bosonic1, bosonic2, bosonic3, bosonic4], and also that Pauli-twirling the amplitude damping channel results in biased 2-Pauli noise [amp_damping_twirl]. Such applications could be combined with our work in the future.
Finally, although noise biases can lead to remarkable threshold improvements, it is well known that such gains can be erased in the experimental setting, principally due to complications with gate operations which can unbias the dominant noise channel [puri2020bias]. As holographic codes approach experimental realization, such considerations will also be necessary.
IV.3 Acknowledgements
@@
We thank Aritra Sarkar, David Elkouss, and Jens Eisert for discussions related to concatenated codes and quantum channel capacity, and Sivaprasad Omanakuttan for mentioning issues related to bias preservation, principally investigated in [puri2020bias]. We also thank Mackenzie Shaw and Kathleen Chang for mentioning at the conference FTQC ’24 in Benasque several realistic examples of 2-Pauli noise arising in practical quantum computing. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation”.
IV.4 Author Contributions